Solve for x and y.
4x+9y=365

Answer by Alan3354(21583)

4x+9y=365
----------
It's not possible. If you have 2 variables, you need 2 equations.
With only 1 eqn, you can find an infinite number of "solutions", eg:
x = 0, y = 365/9
x = 2, y = 119/3
x = 80, y = 5 etc